OpenStudy (anonymous):

What techniques are there to factor quadratic equations that have complex conjugate roots? I do not always want to go through quadratic formula to be able to find these out...

OpenStudy (anonymous):

use the quadratic formula..

OpenStudy (dumbcow):

if the quadratic has complex solutions, then by definition it can't be factored. But if you don't want to use quadratic formula, you can notice the following: Let \[x = a \pm bi\] then \[x-a = \pm bi\] \[(x-a)^{2} = -b^{2}\] \[(x-a)^{2} +b^{2} = 0\] if you can put the quadratic in this form , you will know the complex solutions Note: if b^2 is negative then solutions are real